OpenStudy (anonymous):

Anyone want to help me on Number 2? I have some ideas but I'm not honestly sure...

TheSmartOne (thesmartone):

Sure

OpenStudy (anonymous):

1st year Calculus.

OpenStudy (anonymous):

@nincompoop ,@amistre64

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

solve for R, what do we get?

OpenStudy (anonymous):

it's already solved for R.

OpenStudy (anonymous):

Ohh I see.

OpenStudy (anonymous):

\[R=\frac{ 1 }{ (\frac{ 1 }{ R1 } + \frac{ 1 }{ R2 })}\]

OpenStudy (amistre64):

\[\frac1R=\frac{R_1+R_2}{R_1R_2}\] \[R=\frac{R_1R_2}{R_1+R_2}\]

OpenStudy (anonymous):

yes I did that next.

OpenStudy (anonymous):

I can't see how they got a square root.

OpenStudy (anonymous):

Multiply by conjugate perhaps?

OpenStudy (amistre64):

well, lets find R^2 and see if that helps, then theres the restrictions they provide to test things

OpenStudy (anonymous):

So square both sides?

OpenStudy (amistre64):

if R^2 < R1R2/2 ..... we can start to make somme assumptions

OpenStudy (anonymous):

Yeah but we just have R. I guess we square both sides to make it R^2.

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

this is just an idea at the moment, with any luck it heads us in the right direction

OpenStudy (anonymous):

Yep.

OpenStudy (amistre64):

\[R^2=\frac{(R_1R_2)^2}{R_1^2+R_2^2+2R_1R_2}\]

OpenStudy (anonymous):

Nice latex!

OpenStudy (anonymous):

Yep I got your answer so far...

OpenStudy (amistre64):

now can we compare this with R^2 setup with R1R2/2 to show that its less than it \[\frac{(R_1R_2)^2}{R_1^2+R_2^2+2R_1R_2}<\frac{R_1R_2}{2}\]

OpenStudy (anonymous):

How did you get R1R2/2 ?

OpenStudy (amistre64):

if we cross multiply .... its stated as a criteria already: R < sqrt(R1R2/2) therefore R^2<R1R2/2 for some condition i believe, something greater than 1 maybe \[2(R_1R_2)^2<R_1R_2({R_1^2+R_2^2+2R_1R_2})\]

OpenStudy (anonymous):

Ohh I see. That is what we are trying to show.

OpenStudy (amistre64):

1/3 = sqrt(1/9) 1/9 = 1/9 or maybe no restrictions

OpenStudy (amistre64):

yes, we are trying to determine a reasonable approach, we know R^2 is equal to what we found, now we need to compare it to what is stated

OpenStudy (amistre64):

\[2(R_1R_2)^2<{R_1^3R_2+R_1R_2^3+2(R_1R_2)^2}\] since the R parts are >0 then when we add something to what we have, it gets bigger right?

OpenStudy (anonymous):

But problem. We eliminated the R from the whole equation.

OpenStudy (amistre64):

no we did not eliminate it, we used an equivalent statement for it

OpenStudy (anonymous):

Okay. I just thought we had to match the statement given in the question. But yes this is perfectly valid.

OpenStudy (amistre64):

we said: R = k, then used k as a substutite to work with since it has the same parts that we are trying to proof with

OpenStudy (amistre64):

we could have replaced R by 1/R1 + 1/R2 to start with, but i just took the path i did because im old and stubborn i guess lol

OpenStudy (anonymous):

It's fine :P .

OpenStudy (anonymous):

Divide both sides by R1R2?

OpenStudy (amistre64):

\[\color{red}{R^2}=\left(\frac{(R_1R_2)^2}{R_1^2+R_2^2+2R_1R_2}\right)<\left(\sqrt{\frac{R_1R_2}{2}}\right)^2\]

OpenStudy (amistre64):

1/3 ^2 = 1/9 so when we are less than 1, the square is smaller .....

OpenStudy (amistre64):

this might have to be worked into it somehow

OpenStudy (anonymous):

I'm looking...

OpenStudy (amistre64):

let me gather my idea together to see how valid it is

OpenStudy (anonymous):

Factor the bottom back to what it was before maybe...?

OpenStudy (amistre64):

show that:\[\frac{1}{R_1}+\frac{1}{R_2}<\sqrt{\frac{R_1R_2}{2}}\]

OpenStudy (amistre64):

pfft, sqrt(2/R1R2) since 1/R is = to .... slight typo

OpenStudy (amistre64):

\[\frac{1}{\frac{1}{R_1}+\frac{1}{R_2}}<\sqrt{\frac{R_1R_2}2}\] thats the objective .... agreed?

OpenStudy (anonymous):

I don't see a typo. You're just stating the thing we have to find.

OpenStudy (amistre64):

yeah, just trying to get it subbed in at the start to work ot

OpenStudy (anonymous):

Why the reciprocal?

OpenStudy (anonymous):

Ohh wait nvm. I'm dumb.

OpenStudy (amistre64):

we want to show that R < sqrt(....), but 1/R = 1/R1 + 1/R2

OpenStudy (anonymous):

I get it.

OpenStudy (anonymous):

yeah yeah :P .

OpenStudy (amistre64):

we can take a value in the interval: 0<k<1, say k=1/4 .... sqrt(1/4) = 1/2 or (1/2)^2 < 1/2

OpenStudy (amistre64):

so if R1 and R2 are small .... does this work out well for us

OpenStudy (anonymous):

Define "small" .

OpenStudy (amistre64):

between 0 and 1

OpenStudy (amistre64):

.000000000000000000001 lol

OpenStudy (anonymous):

Okay :P .

OpenStudy (anonymous):

Maybe we can manipulate the LHS like we did before?

OpenStudy (amistre64):

maybe

OpenStudy (anonymous):

I'm trying afew things myself.

OpenStudy (amistre64):

thats always good, since my ideas may or maynot be useful: show that IF \[\frac{1}{\frac{1}{R_1}+\frac{1}{R_2}}<\sqrt{\frac{R_1R_2}2}\] so we assume this to be true 1/3 < 1/2 1/9 < 1/4 still hold true for the squares so my concerns at the moment seem to be without merit, square both sides and the inequality holds

OpenStudy (anonymous):

Induction?

OpenStudy (amistre64):

no need for induction, my concern was for some reason an interval between 0 and 1. had to make sure that i wasnt introducing a bad element

OpenStudy (amistre64):

if 0 < a < b < 1 < c then squaring it all: 0 < a^2 < b^2 < 1 < c^2 the ineuality still holds true

OpenStudy (anonymous):

That is true but how does that help us?

OpenStudy (amistre64):

it simply means that we can work the squares to show that the inequailty is good \[\frac{1}{\frac{1}{R_1}+\frac{1}{R_2}}<\sqrt{\frac{R_1R_2}2}\] \[\frac{R_1R_2}{R_1+R_2}<\sqrt{\frac{R_1R_2}2}\] \[(\frac{R_1R_2}{R_1+R_2})^2<(\sqrt{\frac{R_1R_2}2})^2\] \[\frac{(R_1R_2)^2}{R^2_1+R^2_2+2R_1R_2}<\frac{R_1R_2}2\] since R1 and R2 are positive, then we have no negative stuff to contend with and the sign issues are no problem \[2(R_1R_2)^2<{R_1R_2}({R^2_1+R^2_2+2R_1R_2})\] \[2(R_1R_2)^2<{R^3_1R_2+R_1R^3_2+2(R_1R_2})^2\] subtracting 2(RR)^2 we get \[0<{R^3_1R_2+R_1R^3_2}\] which is true for any R1, R2 > 0 regardless of size

OpenStudy (anonymous):

Wonderful! Thanks :D .

OpenStudy (amistre64):

we can factor out the R1 R2 and divide off the excess to show that 0<R1R2 as well, ut how far you want to take it is up to you

OpenStudy (anonymous):

I'm taking complex analysis and I can't do basic algebra...

OpenStudy (amistre64):

i asked my discrete math teacher: do you know how much algebra i had to forget to learn this?

OpenStudy (anonymous):

Haha xD .